OpenStudy (anonymous):

Find the volume of a cylinder with d = 8.4 m: h = 9.3 m.

OpenStudy (anonymous):

3. Find the volume of a cylinder with d = 8.4 m: h = 9.3 m. A. 55.5896 cubed m. B. 515.123 cubed m C. 67.89 cubed m D. 570.5126 cubed m

OpenStudy (anonymous):

B

OpenStudy (anonymous):

r=1/2 of d.. r=8.4/2=4.2*3.14*9.3=515.123